# LagCeresa.bib

### v1

@@ 3.3. A Lefschetz Jacobian
In this Section, we reformulate the flux maps constructed in 2.1 and 3.2 as “symplectic Abel-Jacobi maps” to a Lefschetz Jacobian. These can be thought of as a symplectic analogue of intermediate Jacobians. We work under the assumption that $X$ is a Lefschetz manifold, that is the maps
$\cup[\omega]^{k}:H^{n-k}(X;\mathbb{R})\longrightarrow H^{n+k}(X;\mathbb{R})$
@@
Letting $\mathcal{L}^{or}$ be a set of suitable oriented Lagrangians in $X$, define $\mathbb{L}^{or}:=\oplus_{L\in\mathcal{L}}\mathbb{Z}\cdot L$. This comes with an obvious cycle class map
$cyc:\mathbb{L}^{or}\longrightarrow H_{n}(X;\mathbb{Z}),$
@@
$\Theta:\hskip 17.07164pt\mathbb{L}^{or}_{hom}\hskip 17.07164pt$ $\longrightarrow LJ(X):=H^{n+1}(X;\mathbb{R})^{\vee}/H_{n+1}(X;\mathbb{Z})$
@@
where $\partial\gamma=L$. This is independent of the choice of such $\gamma$ because we view $H_{n+1}(X;\mathbb{Z})$ as a lattice inside $H^{n+1}(X;\mathbb{R})^{\vee}$ through the usual integration map over $(n+1)$-cycles. To see that $\Theta$ is well-defined at the level of cohomology, notice that an exact $(n+1)$-form $d\nu$ must be of the form $\omega\wedge d\mu=d(\omega\wedge\mu)$ for some $(n-2)$-form $\mu$. Therefore
[…]
Given an algebraic Lagrangian cobordism $\Gamma\subset X\times X(B)$, and given a choice of basepoint $b_{0}\in B$, one can consider the map
$\varphi_{\Gamma}:B$ $\longrightarrow LJ(X)$
@@
It follows from Theorem 3.5 and by the fact that $H^{n+1}(X;\mathbb{R})\cong[\omega]\cup H^{n-1}(X;\mathbb{R})$ that nullcobordant elements in $\mathbb{L}^{or}_{hom}$ lie in the kernel of $\Theta$. In other words, $\varphi_{\Gamma}$ factors as
${B}$${LJ(X)}$${\mathcal{C}ob^{or}(X)_{hom}.}$$\scriptstyle{b\mapsto[\Gamma(F_{b})-\Gamma(F_{b_{0}})]}$$\scriptstyle{\varphi_{\Gamma}}$$\scriptstyle{\tilde{\varphi}_{\Gamma}}$
@@
(3.2) $\left(H^{n+1}(X;\mathbb{R})/([\omega]^{2}\cup H^{n-3}(X;\mathbb{R}))\right)^{\vee}\subset\left(H^{n+1}(X;\mathbb{R})\right)^{\vee}.$
We now describe a tropical structure for $LJ(X)$ by choosing a lattice in $(H^{n+1}(X;\mathbb{R}))^{\vee}$ of those maps $H^{n+1}(X;\mathbb{R})\longrightarrow\mathbb{R}$ which evaluate to integers on a certain sublattice $H^{n+1}(X;\mathbb{R})_{\mathbb{Z}}$ of $H^{n+1}(X;\mathbb{R})$. To define $H^{n+1}(X;\mathbb{R})_{\mathbb{Z}}$, we invoke the Lefschetz decomposition
$H^{n+1}(X;\mathbb{R})\cong\bigoplus_{i}[\omega]^{i}\cup\ker([\omega]^{2i}:H^{n-2i+1}(X;\mathbb{R})\longrightarrow H^{n+2i+1}(X;\mathbb{R})).$
@@
(3.3) $H^{n+1}(X;\mathbb{R})_{\mathbb{Z}}:=\bigoplus_{i}[\omega]^{i}\cup\ker([\omega]^{2i}:H^{n-2i+1}(X;\mathbb{Z})\longrightarrow H^{n+2i+1}(X;\mathbb{R})).$
Unless otherwise stated, from now on when referring to $LJ(X)$ as a tropical object we endow it with the aforementioned tropical structure.
@@
using the observation above. Recall that up to a global coordinate transformation, $B\cong B(M):=\mathbb{R}^{n}/M\cdot\mathbb{Z}^{n}$ for some $M\in GL(n;\mathbb{R})$, where $\mathbb{R}^{n}$ is considered with the tropical structure given by the integer lattice $\mathbb{Z}^{n}\subset\mathbb{R}^{n}$. Given $v\in TB\cong\mathbb{R}^{n}$, there is a corresponding element $[v]\in H_{1}(B;\mathbb{R})$ obtained by identifying $H_{1}(B;\mathbb{R})\cong(M\cdot\mathbb{Z}^{n})\otimes\mathbb{R}\cong TB$. Fixing a point $b\in B$, a fundamental class $[F_{b}]\in H_{n}(F_{b};\mathbb{Z})$, and using the Künneth isomorphism
@@
The fact that the tropical structure defined by 3.3 makes $\varphi_{\Delta}$ into a tropical map follows from the observation that if $v\in T_{\mathbb{Z}}B$, then $[v]\otimes[F_{b}]\in M^{-1}\cdot H_{1}(B;\mathbb{Z})\otimes H_{n}(F_{b};\mathbb{Z}).$ This corresponds to maps $H^{n+1}(X(B);\mathbb{R})\longrightarrow\mathbb{R}$ which evaluate to integers on the lattice
$\psi\left(M^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{n}(F_{b};\mathbb{Z})\right)\subset H^{n+1}(X(B);\mathbb{R}).$
@@
$[\omega]\in\psi\left(M^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{1}(F_{b};\mathbb{Z})\right)$
@@
$M^{-1}\cdot H_{1}(B;\mathbb{Z})\otimes H_{1}(F_{b};\mathbb{Z}).$
@@
$\psi\left(M^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{n}(F_{b};\mathbb{Z})\right)\subset[\omega]\cup\ker([\omega]^{2}:H^{n-1}(X(B);\mathbb{Z})\longrightarrow H^{n+1}(X(B);\mathbb{R})),$
@@
It follows from 3.2 that the proof of tropicality of a map $\varphi_{\Gamma}:B\longrightarrow LJ(X)$ induced by an algebraic Lagrangian cobordism $\Gamma\subset X(B)\times X$ only depends on the choice of tropical structure in first component of the Lefschetz decomposition, that is in
$[\omega]\cup\ker([\omega]^{2}:H^{n-1}(X;\mathbb{R})\longrightarrow H^{n+3}(X;\mathbb{R})).$
Additionally, in this instance where $\Gamma=\Delta\subset X(B)\times X(B)$, the only component relevant to the tropical structure is
$[\omega]\cup\psi\left(H^{0}(B;\mathbb{R})\otimes H^{n-1}(F_{b};\mathbb{R}))\right)\subset[\omega]\cup\ker([\omega]^{2}:H^{n-1}(X;\mathbb{R})\longrightarrow H^{n+3}(X;\mathbb{R})).$
@@
$[\omega]\cup\psi\left(\bigoplus_{i\geq 1}H^{i}(B;\mathbb{R})\otimes H^{n-i-1}(F_{b};\mathbb{R}))\right)$

### v2

@@ 4.3. A Lefschetz Jacobian
In this Section, we reformulate the flux maps constructed in Sections 3.1 and 4.2 as “symplectic Abel-Jacobi maps” to a Lefschetz Jacobian. These can be thought of as a symplectic analogue of intermediate Jacobians. We work under the assumption that $M$ is a Lefschetz manifold, that is the maps
$\cup[\omega]^{k}:H^{n-k}(M;\mathbb{R})\longrightarrow H^{n+k}(M;\mathbb{R})$
@@
Letting $\mathcal{L}^{or}$ be a set of suitable oriented Lagrangians in $M$, define $\mathbb{L}^{or}:=\oplus_{L\in\mathcal{L}}\mathbb{Z}\cdot L$. This comes with an obvious cycle class map
$cyc:\mathbb{L}^{or}\longrightarrow H_{n}(M;\mathbb{Z}),$
@@
$\Theta:\hskip 17.07164pt\mathbb{L}^{or}_{hom}\hskip 17.07164pt$ $\longrightarrow LJ(M):=H^{n+1}(M;\mathbb{R})^{\vee}/H_{n+1}(M;\mathbb{Z})$
@@
where $\partial\gamma=L$. This is independent of the choice of such $\gamma$ because we view $H_{n+1}(M;\mathbb{Z})$ as a lattice inside $H^{n+1}(M;\mathbb{R})^{\vee}$ through the usual integration map over $(n+1)$-cycles. To see that $\Theta$ is well-defined at the level of cohomology, notice that an exact $(n+1)$-form $d\nu$ must be of the form $\omega\wedge d\mu=d(\omega\wedge\mu)$ for some (n2)(n-2)-form $\mu$. Therefore
$\int_{\gamma}d\eta=\int_{L}\omega\wedge\mu=0.$
Given an algebraic Lagrangian cobordism $\Gamma\subset M\times X(B)$, and given a choice of basepoint $b_{0}\in B$, one can consider the map
$\varphi_{\Gamma}:B$ $\longrightarrow LJ(M)$
@@
It follows from Theorem 4.5 and by the fact that $H^{n+1}(M;\mathbb{R})\cong[\omega]\cup H^{n-1}(M;\mathbb{R})$ that nullcobordant elements in $\mathbb{L}^{or}_{hom}$ lie in the kernel of $\Theta$. In other words, $\varphi_{\Gamma}$ factors as
[…]
@@
(4.2) $\left(H^{n+1}(M;\mathbb{R})/([\omega]^{2}\cup H^{n-3}(M;\mathbb{R}))\right)^{\vee}\subset\left(H^{n+1}(M;\mathbb{R})\right)^{\vee}.$
We now describe a tropical structure for $LJ(M)$ by choosing a lattice in $(H^{n+1}(M;\mathbb{R}))^{\vee}$ of those maps $H^{n+1}(M;\mathbb{R})\longrightarrow\mathbb{R}$ which evaluate to integers on a certain sublattice $H^{n+1}(M;\mathbb{R})_{\mathbb{Z}}$ of $H^{n+1}(M;\mathbb{R})$. To define $H^{n+1}(M;\mathbb{R})_{\mathbb{Z}}$, we invoke the Lefschetz decomposition
$H^{n+1}(M;\mathbb{R})\cong\bigoplus_{i}[\omega]^{i}\cup\ker([\omega]^{2i}:H^{n-2i+1}(M;\mathbb{R})\longrightarrow H^{n+2i+1}(M;\mathbb{R})).$
@@
(4.3) $H^{n+1}(M;\mathbb{R})_{\mathbb{Z}}:=\bigoplus_{i}[\omega]^{i}\cup\ker([\omega]^{2i}:H^{n-2i+1}(M;\mathbb{Z})\longrightarrow H^{n+2i+1}(M;\mathbb{R})).$
Unless otherwise stated, from now on when referring to $LJ(M)$ as a tropical object we endow it with the aforementioned tropical structure.
@@
using the observation above. Recall that up to a global coordinate transformation, $B\cong B(Q):=\mathbb{R}^{n}/Q\cdot\mathbb{Z}^{n}$ for some $Q\in GL(n;\mathbb{R})$, where $\mathbb{R}^{n}$ is considered with the tropical structure given by the integer lattice $\mathbb{Z}^{n}\subset\mathbb{R}^{n}$. Given $v\in TB\cong\mathbb{R}^{n}$, there is a corresponding element $v]\in H_{1}(B;\mathbb{R})$ obtained by identifying $H_{1}(B;\mathbb{R})\cong(Q\cdot\mathbb{Z}^{n})\otimes\mathbb{R}\cong TB$. Fixing a point $b\in B$, a fundamental class $F_{b}]\in H_{n}(F_{b};\mathbb{Z})$, and using the Künneth isomorphism
@@
The fact that the tropical structure defined in Equation (4.3) makes $\varphi_{\Delta}$ into a tropical map follows from the observation that if $v\in T_{\mathbb{Z}}B$, then $v]\otimes[F_{b}]\in Q^{-1}\cdot H_{1}(B;\mathbb{Z})\otimes H_{n}(F_{b};\mathbb{Z}).$ This corresponds to maps $H^{n+1}(X(B);\mathbb{R})\longrightarrow\mathbb{R}$ which evaluate to integers on the lattice
$\psi\left(Q^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{n}(F_{b};\mathbb{Z})\right)\subset H^{n+1}(X(B);\mathbb{R}).$
@@
$\omega]\in\psi\left(Q^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{1}(F_{b};\mathbb{Z})\right)$
@@
$Q^{-1}\cdot H_{1}(B;\mathbb{Z})\otimes H_{1}(F_{b};\mathbb{Z}).$
@@
$\psi\left(Q^{t}\cdot H^{1}(B;\mathbb{Z})\otimes H^{n}(F_{b};\mathbb{Z})\right)\subset[\omega]\cup\ker([\omega]^{2}:H^{n-1}(X(B);\mathbb{Z})\longrightarrow H^{n+1}(X(B);\mathbb{R})),$
@@
It follows from Equation (4.2) that the proof of tropicality of a map $\varphi_{\Gamma}:B\longrightarrow LJ(M)$ induced by an algebraic Lagrangian cobordism $\Gamma\subset\overline{X(B)}\times M$ only depends on the choice of tropical structure in first component of the Lefschetz decomposition, that is in
$\omega]\cup\ker([\omega]^{2}:H^{n-1}(M;\mathbb{R})\longrightarrow H^{n+3}(M;\mathbb{R})).$
Additionally, in this instance where $\Gamma=\Delta\subset\overline{X(B)}\times X(B)$, the only component relevant to the tropical structure is
$\omega]\cup\psi\left(H^{0}(B;\mathbb{R})\otimes H^{n-1}(F_{b};\mathbb{R}))\right)\subset[\omega]\cup\ker([\omega]^{2}:H^{n-1}(M;\mathbb{R})\longrightarrow H^{n+3}(M;\mathbb{R})).$
@@
$\omega]\cup\psi\left(\bigoplus_{i\geq 1}H^{i}(B;\mathbb{R})\otimes H^{n-i-1}(F_{b};\mathbb{R}))\right)$